# Low-degree mod 2 cohomology of classifying spaces of $G_{2}$-gauge groups

### v2

@@ Abstract.
Let $G$ be a simply connected compact simple Lie group and let $\mathcal{G}_{k}$ denote the gauge group of a principal GG–bundle over $S^{4}$ with second Chern class $k\in\pi_{4}(BG)\cong\mathbb{Z}$. For $G=G_{2}$, the pp–local homotopy types of the gauge groups have been completely classified by Kishimoto–Theriault–Tsutaya and Kameko in terms of the order of the fundamental Samelson product $\langle i_{3},1\rangle\in[\Sigma^{3}G_{2},G_{2}]$.
@@
together with Serre and Eilenberg–Moore spectral sequences, we study the Serre spectral sequence
@@
$H^{j}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})=0\quad\text{for }1\leq j\leq 4,\qquad H^{5}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})\neq 0,$
so the first positive-degree generator of the fibre cohomology occurs in degree $5$. As a consequence, there is a distinguished class
@@
whose only possible Serre differential in total degree $\leq 10$ is a $d_{6}$–differential
$d_{6}(u_{5})=\varepsilon(k)\,x_{6}$
from $u_{5}$ to the degree-$6$ generator $x_{6}\in H^{6}(BG_{2};\mathbb{F}_{2})$, for a scalar $\varepsilon(k)\in\mathbb{F}_{2}$ encoding the low-degree effect of the bundle class. In addition, 22–locally we prove that $\varepsilon(k)$ is 44–periodic in $k$ (i.e. it depends only on $k\bmod 4$) and that $\varepsilon(k)=0$ for all $k\equiv 0\pmod{4}$.
@@ 1.1. Gauge groups and their homotopy types
The case $G=G_{2}$ is particularly interesting because of the exceptional nature of $G_{2}$. Kishimoto, Theriault and Tsutaya proved that the classification for $G_{2}$–gauge groups reduces to determining the order of the Samelson product
@@
where $i_{3}\colon S^{3}\to G_{2}$ is the inclusion of the bottom cell and $1$ denotes the identity map [8].
Very recently, Kameko [7] resolved this issue at the prime $2$ and completed the pp–local classification of $G_{2}$–gauge groups. He proved that the order of $\langle i_{3},1\rangle$ is exactly $84$ [7, Theorem 1.1]. Consequently, the pp–local homotopy type of $\mathcal{G}_{k}(G_{2})$ depends only on the greatest common divisor $(k,84)$. Thus the homotopy types of the gauge groups are completely understood up to pp–localization. It is natural to ask how far this information is reflected in the (co)homology of their classifying spaces.
@@ 1.2. Cohomology of $B\mathcal{G}_{k}$ and the Steenrod algebra
Gottlieb identified the classifying space of $\mathcal{G}_{k}$ with a component of a mapping space [5]: specifically, $B\mathcal{G}_{k}$ is homotopy equivalent to $\mathrm{Map}_{k}(S^{4},BG)$, the connected component of the mapping space $\mathrm{Map}(S^{4},BG)$ containing the classifying map $k\colon S^{4}\to BG$. This identification yields a fundamental evaluation fibration
@@
By contrast, there seems to be no published description of the full mod $2$ cohomology ring $H^{*}(B\mathcal{G}_{k};\mathbb{F}_{2})$ for any nontrivial $k$, nor of its structure as an unstable $\mathcal{A}$–module. Our aim here is to give a precise structural description of $H^{*}(B\mathcal{G}_{k};\mathbb{F}_{2})$ in low degrees, in terms of the evaluation fibration and the first nontrivial cohomology of the triple loop space $\Omega^{3}_{0}G_{2}$.
@@ 1.3. Main result: a low-degree description
which is the triple adjoint of the Samelson product $\langle k\cdot i_{3},1\rangle$ by work of Kishimoto–Theriault–Tsutaya [8]. Kameko’s calculation of the 22–primary order of $\langle i_{3},1\rangle$ implies that, after localization at $2$, the map $\partial_{k}$ depends only on $k\bmod 4$.
@@
    $E_{2}^{*,*}(k)^{\leq 10}\;\cong\;H^{*}(BG_{2})\otimes M^{\leq 10},$
    with $E_{2}^{s,t}=H^{s}(BG_{2})\otimes M^{t}$.
@@
    Using the low-degree homotopy groups of $G_{2}$ and the Hurewicz theorem, one has
    $M^{j}=H^{j}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})=0\quad\text{for }1\leq j\leq 4,\qquad M^{5}\neq 0.$
    In particular, there exists a nonzero class
@@
    and $u_{5}$ is the first positive-degree generator of $H^{*}(\Omega^{3}_{0}G_{2})$.
@@
    In total degree $j\leq 10$ the differentials
    $d_{r}\colon E_{r}^{0,t}(k)\longrightarrow E_{r}^{r,t-r+1}(k)$
    with $2\leq r\leq 5$ vanish on all classes in bidegree $(0,t)$, and the only page on which a differential with source in $E_{r}^{0,5}(k)$ can be nonzero is $r=6$. After choosing a generator $x_{6}\in H^{6}(BG_{2})\cong E_{2}^{6,0}(k)$ and $u_{5}\in M^{5}\cong E_{2}^{0,5}(k)$ there is a uniquely determined scalar $\varepsilon(k)\in\mathbb{F}_{2}$ such that
@@
    For each $j\leq 10$ the spectral sequence induces a finite filtration of $H^{j}(B\mathcal{G}_{k})$ whose associated graded contains, as a direct summand, the image of $U$ in $E_{\infty}^{*,*}(k)^{j}$. This summand is explicitly computable from $H^{*}(BG_{2})\otimes M^{\leq 10}$ together with $\varepsilon(k)$, while the remaining extension term depends only on $M^{\leq j}$ and the multiplicative structure of the Serre spectral sequence.
@@
Theorem 1.1 should be viewed as a structural reduction. It does not claim a closed formula for $\varepsilon(k)$, nor does it assert that $\varepsilon(k)$ is nontrivial. Kameko’s result shows that the 22–local homotopy type of $\mathcal{G}_{k}(G_{2})$ depends on $k\bmod 4$ through the 22–primary order of the Samelson product $\langle i_{3},1\rangle$ [8, 7]. It is therefore natural to expect that $\varepsilon(k)$ is governed by the same 22–primary data.
In fact, we prove (Lemma 4.5) that, 22–locally, the scalar $\varepsilon(k)$ depends only on the residue class of $k$ modulo $4$, and moreover $\varepsilon(k)=0$ whenever $4\mid k$ (Corollary 4.6). Thus the remaining low-degree problem reduces to determining the two values $\varepsilon(1)$ and $\varepsilon(2)$.
@@
The paper is organised as follows. In Section 2 we review the necessary background on gauge groups, Samelson products, the mod $2$ cohomology of $BG_{2}$ and the Steenrod algebra. In Section 3 we describe the evaluation fibration and its Serre spectral sequence, and we recall the relationship between the connecting map and the Samelson product. Finally, Section 4 contains the proof of Theorem 1.1, including a detailed analysis of the bidegrees for $j\leq 10$ and the vanishing of $H^{j}(\Omega^{3}_{0}G_{2})$ for $1\leq j\leq 4$.
@@ 2.1. Gauge groups and mapping spaces
Following the notation of [8, 7], we write $\mathcal{G}_{k}=\mathcal{G}_{k}(G_{2})$ for the gauge group of the principal $G_{2}$–bundle classified by $k\in\pi_{4}(BG_{2})\cong\mathbb{Z}$. By a classical result of Gottlieb [5], the classifying space $B\mathcal{G}_{k}$ is homotopy equivalent to $\mathrm{Map}_{k}(S^{4},BG_{2})$, the connected component of the mapping space $\mathrm{Map}(S^{4},BG_{2})$ containing maps of degree $k$.
@@
where the fibre $\mathrm{Map}_{*}(S^{4},BG_{2})_{k}$ is homotopy equivalent to the triple loop space $\Omega^{3}_{0}G_{2}$.
@@ 2.2. Samelson products and the connecting map
where $i_{3}\colon S^{3}\to G$ is the inclusion of the bottom cell and $1$ is the identity. Kishimoto, Theriault and Tsutaya showed that the connecting map
@@
is the triple adjoint of the Samelson product $\langle k\cdot i_{3},1\rangle$ [8, Lemma 2.1]. In particular $\partial_{k}\simeq k\cdot\partial_{1}$, and the order of $\partial_{1}$ is equal to the order of $\langle i_{3},1\rangle$.
For $G=G_{2}$, the work of Kishimoto–Theriault–Tsutaya [8] and Kameko [7] implies that the 22–primary component of $\langle i_{3},1\rangle$ has order $4$. Thus, after localizing at $2$, the homotopy class of $\partial_{k}$ depends only on $k\bmod 4$. This 22–primary information is the only input from the classification of $G_{2}$–gauge groups that will enter our cohomological analysis.
@@ 3.2. The triple loop space $\Omega^{3}_{0}G_{2}$
    The first positive-degree cohomology group $M^{5}$ is nonzero:
    $M^{5}=H^{5}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})\neq 0.$
    In particular, there exists a nonzero class $u_{5}\in M^{5}$.
@@
We work integrally and then pass to mod $2$ coefficients.
Step 1: Low-degree homology of $\Omega^{3}_{0}G_{2}$. By Lemma 3.1, the space $X=\Omega^{3}_{0}G_{2}$ is 22–connected. The Hurewicz theorem for (n1)(n-1)–connected spaces states that if $X$ is (n1)(n-1)–connected with $n\geq 2$, then $H_{i}(X;\mathbb{Z})=0$ for $1\leq i\leq n-1$, and the Hurewicz map
$h\colon\pi_{n}(X)\longrightarrow H_{n}(X;\mathbb{Z})$
is an isomorphism (see, for example, [6, Theorem 4.32]). In our case $n=3$, so $X$ is 22–connected and we obtain:
$H_{1}(X;\mathbb{Z})=H_{2}(X;\mathbb{Z})=0,\qquad h\colon\pi_{i}(X)\xrightarrow{\cong}H_{i}(X;\mathbb{Z})\quad\text{for }i=3,4,5.$
Using $\pi_{i}(X)\cong\pi_{i+3}(G_{2})$ for $i\geq 0$ and the low-degree homotopy groups of $G_{2}$ quoted above, we have
$H_{3}(X;\mathbb{Z})$ $\cong\pi_{3}(X)\cong\pi_{6}(G_{2})\quad\text{(a finite $3$--group)},$
$H_{4}(X;\mathbb{Z})$ $\cong\pi_{4}(X)\cong\pi_{7}(G_{2})=0,$
$H_{5}(X;\mathbb{Z})$ $\cong\pi_{5}(X)\cong\pi_{8}(G_{2}),$
and $\pi_{8}(G_{2})$ contains 22–torsion by the computations of Mimura and Toda [11]. Thus $H_{5}(X;\mathbb{Z})$ has nontrivial 22–primary torsion.
Step 2: Low-degree cohomology with $\mathbb{F}_{2}$–coefficients. We now pass to mod $2$ cohomology via the universal coefficient short exact sequence
$0\longrightarrow\operatorname{Ext}^{1}_{\mathbb{Z}}(H_{i-1}(X;\mathbb{Z}),\mathbb{F}_{2})\longrightarrow H^{i}(X;\mathbb{F}_{2})\longrightarrow\operatorname{Hom}_{\mathbb{Z}}(H_{i}(X;\mathbb{Z}),\mathbb{F}_{2})\longrightarrow 0.$
For $i=1$ we use $H_{0}(X;\mathbb{Z})\cong\mathbb{Z}$ and $H_{1}(X;\mathbb{Z})=0$ to obtain $H^{1}(X;\mathbb{F}_{2})=0$. For $i=2$ we use $H_{1}(X;\mathbb{Z})=H_{2}(X;\mathbb{Z})=0$ and obtain $H^{2}(X;\mathbb{F}_{2})=0$. For $i=3$ we have $H_{2}(X;\mathbb{Z})=0$, so
$H^{3}(X;\mathbb{F}_{2})\;\cong\;\operatorname{Hom}_{\mathbb{Z}}(H_{3}(X;\mathbb{Z}),\mathbb{F}_{2}).$
Since $H_{3}(X;\mathbb{Z})\cong\pi_{6}(G_{2})$ is a finite 33–group, and $3$ is invertible in $\mathbb{F}_{2}$, any homomorphism $H_{3}(X;\mathbb{Z})\to\mathbb{F}_{2}$ must be trivial, hence $H^{3}(X;\mathbb{F}_{2})=0$.
For $i=4$ we have $H_{3}(X;\mathbb{Z})$ a 33–group and $H_{4}(X;\mathbb{Z})=0$, so
$H^{4}(X;\mathbb{F}_{2})\;\cong\;\operatorname{Ext}^{1}_{\mathbb{Z}}(H_{3}(X;\mathbb{Z}),\mathbb{F}_{2}).$
For a cyclic group $\mathbb{Z}/n$ one has
$\operatorname{Ext}^{1}_{\mathbb{Z}}(\mathbb{Z}/n,\mathbb{F}_{2})\cong\mathbb{F}_{2}/n\mathbb{F}_{2},$
which is zero whenever $n$ is odd. Since $H_{3}(X;\mathbb{Z})$ is a 33–group, it follows that $\operatorname{Ext}^{1}_{\mathbb{Z}}(H_{3}(X;\mathbb{Z}),\mathbb{F}_{2})=0$, hence $H^{4}(X;\mathbb{F}_{2})=0$.
This proves $M^{1}=M^{2}=M^{3}=M^{4}=0$; the equality $M^{0}\cong\mathbb{F}_{2}$ follows from path-connectedness. This establishes (a).
For $i=5$ we use the fact that $H_{4}(X;\mathbb{Z})=0$, so
$H^{5}(X;\mathbb{F}_{2})\;\cong\;\operatorname{Hom}_{\mathbb{Z}}(H_{5}(X;\mathbb{Z}),\mathbb{F}_{2}).$
Since $H_{5}(X;\mathbb{Z})\cong\pi_{8}(G_{2})$ contains 22–torsion, there is a nonzero homomorphism $H_{5}(X;\mathbb{Z})\to\mathbb{F}_{2}$, and hence $H^{5}(X;\mathbb{F}_{2})\neq 0$. This proves (b). ∎
For the remainder of the paper we fix a nonzero element $u_{5}\in M^{5}=H^{5}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$.

### v4

@@ Abstract.
Let $G$ be a simply connected compact simple Lie group and let $\mathcal{G}_{k}$ denote the gauge group of a principal GG–bundle over $S^{4}$ with second Chern class $k\in\pi_{4}(BG)\cong\mathbb{Z}$. For $G=G_{2}$, the pp–local homotopy types of the gauge groups have been completely classified through the work of Kishimoto–Theriault–Tsutaya and Kameko, in terms of the order of the fundamental Samelson product $\langle i_{3},1\rangle\in[\Sigma^{3}G_{2},G_{2}]$.
@@
and the associated Serre spectral sequence, together with low-degree information on $\Omega^{3}_{0}G_{2}$, we study
@@
$H^{j}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})=0\quad\text{for }1\leq j\leq 4,\qquad H^{5}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})\cong\mathbb{F}_{2},$
so the least positive degree in which the fibre cohomology is nonzero is $5$. As a consequence, there is a distinguished generator
@@
such that, in total degree $\leq 10$, the only possible Serre differential with source $u_{5}$ is a $d_{6}$–differential
$d_{6}(u_{5})=\varepsilon(k)\,x_{6},$
where $x_{6}\in H^{6}(BG_{2};\mathbb{F}_{2})$ is the degree-$6$ generator and $\varepsilon(k)\in\mathbb{F}_{2}$ records the contribution of the bundle class to this first nontrivial fibre differential. In addition, 22–locally we prove that $\varepsilon(k)$ is 88–periodic in $k$ (i.e. it depends only on $k\bmod 8$) and that $\varepsilon(k)=0$ for all $k\equiv 0\pmod{8}$.
@@
Acknowledgment
The author would like to express his sincere gratitude to the anonymous referee for the careful reading of the manuscript and for the valuable suggestions and corrections, which have greatly improved the clarity and overall quality of the paper.
@@ 1.1. Gauge groups and their homotopy types
The case $G=G_{2}$ is particularly interesting because of the exceptional nature of $G_{2}$. Kishimoto, Theriault and Tsutaya  [11] proved that the classification for $G_{2}$–gauge groups reduces to determining the order of the Samelson product
@@
where $i_{3}\colon S^{3}\to G_{2}$ is the inclusion of the bottom cell and $1$ denotes the identity map. Moreover, their work determines the number of homotopy types up to one factor of $2$ [11]. Kameko subsequently resolved the remaining 22–primary ambiguity by proving that the order of $\langle i_{3},1\rangle$ is exactly $168$ [10, Theorem 1.4]. Consequently, the pp–local homotopy classification of $\mathcal{G}_{k}(G_{2})$ is determined by the greatest common divisor $(k,168)$; at the prime $2$, it is determined by $(k,8)$ [10, Theorem 1.1 and Proposition 1.2]. Thus the homotopy types of the gauge groups are completely understood up to pp–localization. It is natural to ask how far this information is reflected in the (co)homology of their classifying spaces.
@@ 1.2. Cohomology of $B\mathcal{G}_{k}$ and the Steenrod algebra
Gottlieb identified the classifying space of $\mathcal{G}_{k}$ with a component of a mapping space [7]: specifically, $B\mathcal{G}_{k}$ is homotopy equivalent to $\mathrm{Map}_{k}(S^{4},BG)$, the path component of $\mathrm{Map}(S^{4},BG)$ containing a classifying map representing $k\in\pi_{4}(BG)\cong\mathbb{Z}$. This identification yields a fundamental evaluation fibration
@@
By contrast, there seems to be no published description of the full mod $2$ cohomology ring $H^{*}(B\mathcal{G}_{k};\mathbb{F}_{2})$ for nontrivial $G_{2}$–bundles over $S^{4}$, nor of its structure as an unstable $\mathcal{A}$–module. There are, however, important partial precedents. Choi studied the mod $p$ homology of classifying spaces of $\mathrm{Sp}(n)$–gauge groups over $S^{4}$ [4], and he also computed the homology of the $G_{2}$–gauge groups themselves using Eilenberg–Moore and Serre spectral sequences [5]. Our aim here is to give a precise structural description of $H^{*}(B\mathcal{G}_{k};\mathbb{F}_{2})$ in low degrees, in terms of the evaluation fibration and the first nontrivial cohomology of the triple loop space $\Omega^{3}_{0}G_{2}$.
@@ 1.3. Main result: a low-degree description
which, by Lang’s description of the boundary map of the evaluation fibration and its formulation in the gauge-group setting [11, 13], is the triple adjoint of the Samelson product $\langle k\cdot i_{3},1\rangle$. Kameko’s calculation that the 22–primary order of $\langle i_{3},1\rangle$ is $8$ implies that, after localization at $2$, the map $\partial_{k}$ depends only on $k\bmod 8$ [10, Proposition 1.3].
@@
    $E_{2}^{*,*}(k)^{\leq 10}\;\cong\;\bigl(H^{*}(BG_{2})\otimes M\bigr)^{\leq 10},$
    where the superscript ≤10 denotes truncation in total degree. Equivalently, $E_{2}^{s,t}(k)\cong H^{s}(BG_{2})\otimes M^{t}$ for $s+t\leq 10$.
@@
    Using the low-degree homotopy groups of $G_{2}$ together with integral and 22–local Hurewicz arguments, one has
    $M^{j}=H^{j}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})=0\quad\text{for }1\leq j\leq 4,\qquad M^{5}\cong\mathbb{F}_{2}.$
    In particular, there exists a generator
@@
    and $u_{5}$ is a generator of least positive degree in $H^{*}(\Omega^{3}_{0}G_{2})$.
@@
    In total degree $j\leq 10$, all differentials
    $d_{r}(u_{5})\colon E_{r}^{0,5}(k)\longrightarrow E_{r}^{r,\,5-r+1}(k)$
    vanish for $2\leq r\leq 5$ and for $r\geq 7$, and the only page on which $u_{5}$ can support a nontrivial differential is $r=6$. After choosing a generator $x_{6}\in H^{6}(BG_{2})\cong E_{2}^{6,0}(k)$ and the generator $u_{5}\in M^{5}\cong E_{2}^{0,5}(k)$, there is a uniquely determined scalar $\varepsilon(k)\in\mathbb{F}_{2}$ such that
@@
    For each $j\leq 10$, the spectral sequence induces a finite filtration
    $0=F^{j+1}H^{j}(B\mathcal{G}_{k})\subseteq F^{j}H^{j}(B\mathcal{G}_{k})\subseteq\cdots\subseteq F^{0}H^{j}(B\mathcal{G}_{k})=H^{j}(B\mathcal{G}_{k})$
    whose associated graded object is
    $\mathrm{gr}\,H^{j}(B\mathcal{G}_{k})\cong\bigoplus_{s+t=j}E_{\infty}^{s,t}(k).$
    In total degree $j\leq 10$, the contribution to $E_{\infty}^{*,*}(k)$ coming from the $H^{*}(BG_{2})$–submodule generated by $u_{5}$ is explicitly determined by the single scalar $\varepsilon(k)\in\mathbb{F}_{2}$.
@@
Theorem 1.1 should be viewed as a structural reduction. It does not claim a closed formula for $\varepsilon(k)$, nor does it assert that $\varepsilon(k)$ is nontrivial. Kameko’s calculation shows that the 22–primary order of $\langle i_{3},1\rangle$ is $8$, while the 22–local homotopy classification of the gauge groups is indexed by $(k,8)$ [10, Proposition 1.2 and Proposition 1.3]. It is therefore natural to expect that $\varepsilon(k)$ is governed by the same 22–primary data.
In fact, we prove (Lemma 4.4) that, 22–locally, the scalar $\varepsilon(k)$ depends only on the residue class of $k$ modulo $8$, and moreover $\varepsilon(k)=0$ whenever $8\mid k$ (Corollary 4.5). Thus the argument given here reduces the remaining low-degree problem to determining the seven values $\varepsilon(1),\ldots,\varepsilon(7)$. Although Kameko’s classification identifies the 22–local homotopy types of the gauge groups by the invariant $(k,8)$, the proof below uses the stronger condition that the connecting maps agree, namely congruence modulo $8$. Any further reduction among these residue classes would require an additional fibrewise comparison argument, which we do not establish here.
@@
Remark 1.3 (Applications and significance).
Although Theorem 1.1 is a low-degree result, it has several concrete applications.
First, it provides a first cohomological manifestation of the dependence of the classifying spaces $B\mathcal{G}_{k}$ on the bundle class $k$, through the first nontrivial fibre differential. In this sense, the scalar $\varepsilon(k)\in\mathbb{F}_{2}$ may be viewed as a low-degree cohomological shadow of the 22–local homotopy classification of the gauge groups $\mathcal{G}_{k}(G_{2})$.
Second, the theorem reduces the first kk–dependent part of the low-degree analysis of $H^{*}(B\mathcal{G}_{k};\mathbb{F}_{2})$ to two more concrete problems: determining the next stages of the fibre cohomology $H^{*}(\Omega_{0}^{3}G_{2};\mathbb{F}_{2})$ beyond degree $5$, and determining the values of the scalar $\varepsilon(k)$. Thus the remaining difficulty becomes much more explicit and the theorem provides a practical framework for further calculations.
Third, since our goal is to study $H^{*}(B\mathcal{G}_{k};\mathbb{F}_{2})$ as an unstable module over the Steenrod algebra, the present low-degree description gives the first structural input for any future analysis of generators, Steenrod operations, and hit-theoretic questions for these classifying spaces.
Finally, because $\mathcal{G}_{k}\simeq\Omega B\mathcal{G}_{k}$, the low-degree information obtained here is also relevant to comparisons with the mod $2$ homology of the gauge groups themselves, and may serve as a starting point for further work relating the classifying-space approach of the present paper to the Eilenberg–Moore and Serre spectral sequence calculations of Choi [5].
In particular, the theorem isolates the first possible obstruction to the survival of the degree-$6$ base class in the evaluation Serre spectral sequence, and thus identifies the first place where the bundle class can affect the mod $2$ cohomology of $B\mathcal{G}_{k}$ through the first nontrivial fibre class.
The paper is organised as follows. In Section 2, we review the necessary background on gauge groups, Samelson products, the mod $2$ cohomology of $BG_{2}$ and the Steenrod algebra. In Section 3, we describe the evaluation fibration and its Serre spectral sequence, and we recall the relationship between the connecting map and the Samelson product. Finally, Section 4 contains the proof of Theorem 1.1, including a detailed analysis of the bidegrees for $j\leq 10$ and the vanishing of $H^{j}(\Omega^{3}_{0}G_{2})$ for $1\leq j\leq 4$.
@@ 2.1. Gauge groups and mapping spaces
Following the notation of [10, 11], we write $\mathcal{G}_{k}=\mathcal{G}_{k}(G_{2})$ for the gauge group of the principal $G_{2}$–bundle classified by $k\in\pi_{4}(BG_{2})\cong\mathbb{Z}$. By a classical result of Gottlieb [7], the classifying space $B\mathcal{G}_{k}$ is homotopy equivalent to $\mathrm{Map}_{k}(S^{4},BG_{2})$, the connected component of the mapping space $\mathrm{Map}(S^{4},BG_{2})$ corresponding to the class $k\in\pi_{4}(BG_{2})\cong\mathbb{Z}$.
@@
where $\mathrm{Map}_{*}(S^{4},BG_{2})_{k}$ denotes the path component of the based mapping space $\mathrm{Map}_{*}(S^{4},BG_{2})$ corresponding to $k\in\pi_{4}(BG_{2})\cong\mathbb{Z}$. For simply connected $G$, this fibre is homotopy equivalent to $\Omega_{0}^{3}G$.
@@ 2.2. Samelson products and the connecting map
where $i_{3}\colon S^{3}\to G$ is the inclusion of the bottom cell and $1$ is the identity. Lang [13] showed that the boundary map of the evaluation fibration is identified with the adjoint of a Whitehead product. In the present gauge-group setting, this implies that the connecting map
@@
is the triple adjoint of the Samelson product $\langle k\cdot i_{3},1\rangle$; see [11, Lemma 2.1] for the formulation used here. In particular $\partial_{k}\simeq k\cdot\partial_{1}$, and the order of $\partial_{1}$ is equal to the order of $\langle i_{3},1\rangle$.
For $G=G_{2}$, Kameko’s corrected calculation [10, Proposition 1.3] implies that the 22–primary component of $\langle i_{3},1\rangle$ has order $8$. Thus, after localizing at $2$, the homotopy class of $\partial_{k}$ depends only on $k\bmod 8$. This 22–primary information is the only input from the classification of $G_{2}$–gauge groups that will enter our cohomological analysis.
@@ 3.2. The triple loop space $\Omega^{3}_{0}G_{2}$
    One has
    $M^{5}=H^{5}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})\cong\mathbb{F}_{2}.$
    In particular, if $0\neq u_{5}\in M^{5}$, then $u_{5}$ is a generator of $M^{5}$ and is of least positive degree in $H^{*}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$.
@@
Let $X=\Omega^{3}_{0}G_{2}$. By Lemma 3.1, $X$ is 22–connected.
Step 1: integral information in degrees $\leq 4$. Since $X$ is 22–connected, the Hurewicz theorem gives
$H_{1}(X;\mathbb{Z})=H_{2}(X;\mathbb{Z})=0,$
an isomorphism
$h\colon\pi_{3}(X)\xrightarrow{\cong}H_{3}(X;\mathbb{Z}),$
and a surjection
[…]
Using $\pi_{i}(X)\cong\pi_{i+3}(G_{2})$ for $i\geq 1$ and the low-degree homotopy groups of $G_{2}$, we obtain
$\pi_{3}(X)\cong\pi_{6}(G_{2})\quad\text{(a finite $3$--group)},\qquad\pi_{4}(X)\cong\pi_{7}(G_{2})=0.$
Hence
$H_{3}(X;\mathbb{Z})\cong\pi_{3}(X)\cong\pi_{6}(G_{2})$
is a finite 33–group, while
$H_{4}(X;\mathbb{Z})=0.$
Now apply the universal coefficient short exact sequence for cohomology. Since $H_{1}(X;\mathbb{Z})=H_{2}(X;\mathbb{Z})=0$, we have
$H^{1}(X;\mathbb{F}_{2})=H^{2}(X;\mathbb{F}_{2})=0.$
Since $H_{3}(X;\mathbb{Z})$ is a finite 33–group, both
$\operatorname{Hom}_{\mathbb{Z}}(H_{3}(X;\mathbb{Z}),\mathbb{F}_{2})=0\qquad\text{and}\qquad\operatorname{Ext}^{1}_{\mathbb{Z}}(H_{3}(X;\mathbb{Z}),\mathbb{F}_{2})=0.$
Using also $H_{4}(X;\mathbb{Z})=0$, it follows that
$H^{3}(X;\mathbb{F}_{2})=H^{4}(X;\mathbb{F}_{2})=0.$
This proves part (a).
Step 2: the first nonzero class in degree $5$. We now work 22–locally. Since $X$ is nilpotent, localization at $2$ induces an isomorphism on mod $2$ homology
[…]
Moreover, for each $i\geq 1$ one has
$\pi_{i}(X_{(2)})\cong\pi_{i}(X)\otimes\mathbb{Z}_{(2)}.$
Because
$\pi_{1}(X)=\pi_{2}(X)=0,\qquad\pi_{3}(X)\cong\pi_{6}(G_{2})\ \text{is $3$--primary},\qquad\pi_{4}(X)\cong\pi_{7}(G_{2})=0,$
it follows that
$\pi_{i}(X_{(2)})=0\qquad\text{for }1\leq i\leq 4.$
Thus $X_{(2)}$ is 44–connected.
Applying the integral Hurewicz theorem to the 44–connected space $X_{(2)}$, we obtain an isomorphism
[…]
Using again $\pi_{i}(X)\cong\pi_{i+3}(G_{2})$ together with the classical computation
$\pi_{8}(G_{2})\cong\mathbb{Z}/2,$
we get
$\pi_{5}(X_{(2)})\cong\pi_{5}(X)\otimes\mathbb{Z}_{(2)}\cong\pi_{8}(G_{2})\otimes\mathbb{Z}_{(2)}\cong\mathbb{Z}/2.$
Hence
$H_{5}(X_{(2)};\mathbb{Z})\cong\mathbb{Z}/2.$
Reducing coefficients mod $2$ gives
$H_{5}(X_{(2)};\mathbb{F}_{2})\cong\mathbb{F}_{2}.$
Using invariance of mod $2$ homology under 22–localization, we conclude that
$H_{5}(X;\mathbb{F}_{2})\cong\mathbb{F}_{2}.$
Since coefficients are taken in the field $\mathbb{F}_{2}$, the universal coefficient theorem gives
$H^{5}(X;\mathbb{F}_{2})\cong\operatorname{Hom}_{\mathbb{F}_{2}}(H_{5}(X;\mathbb{F}_{2}),\mathbb{F}_{2})\cong\mathbb{F}_{2}.$
This proves part (b). ∎
For the remainder of the paper, we fix a generator
$u_{5}\in M^{5}=H^{5}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})\cong\mathbb{F}_{2}.$